OpenStudy (anonymous):

Choose the correct simplification of the expression (8x^2)^3.

OpenStudy (anonymous):

8x^6

OpenStudy (anonymous):

Did that help?

OpenStudy (anonymous):

Because that is a Power of a product so mulitply the exponents together.

OpenStudy (anonymous):

No, you forgot to cube the 8.

OpenStudy (anonymous):

Yea man thank you i got it did you see my other question? haha i @'d you and everything @JasonArnold

OpenStudy (anonymous):

(8x^2)^3 = (8^3)(x^2)^3 = 512x^6

OpenStudy (anonymous):

My Openstudy has a bug like my messages and stuff dont pop up unless I log out then login sorry man!! :D